# Scalability of Metropolis-within-Gibbs schemes for high-dimensional Bayesian models

### v1

@@ A.2 Convergence of the stationary distributions does not imply convergence of the conductances
Appendix B Alternative definition of $s$-conductance
@@
Note that $\Phi_{s}(P)\geq\tilde{\Phi}_{s}(P)$ for every $s\geq 0$ and $\tilde{\Phi}_{0}(P)=\Phi_{0}(P)=\Phi(P)$.
@@
Let $P$ be a coordinate-wise operator as in (5), with target distribution $\pi\in\mathcal{P}(\mathcal{X})$, and $\kappa(P,K)$ as in (7). Then for every $20\leq s\leq 1/$ 2 and $K\subseteq\mathcal{X}$ we have
@@
$P(\partial A)$ $=\frac{1}{d}\sum_{i=1}^{d}P_{i}(\partial A)\geq\frac{1}{d}\sum_{i=1}^{d}\kappa_{i}(P_{i},\mathcal{X})G_{i}(\partial A)\geq\kappa(P)\frac{1}{d}\sum_{i=1}^{d}G_{i}(\partial A).$
@@
Appendix C Regularity assumptions (B4)-(B6) for model (20)
@@
with $s,s^{\prime}=1,\dots S$ and $d=1,\dots,D$. Moreover we write $B_{\delta}$ for the ball of center $\psi^{*}$ and radius $\delta$, and denote expectations with respect to the law of $Y_{j}$ as defined in $(B1)$ by $E_{Y_{j}}[\cdot]$. Finally, we define the posterior characteristic function of $T(\theta_{j})=\left(T_{1}(\theta_{j}),\dots,T_{S}(\theta_{j})\right)$ and $\sum_{j=1}^{k}T(\theta_{j})$, given $\psi$, as $\varphi\left(t\mid Y_{j},\psi\right)=E\left[e^{it^{\top}T(\theta_{j})}\mid Y_{j},\psi\right]$ for $t\in\mathbb{R}^{S}$. and $\varphi^{(k)}\left(t\mid Y_{1:k},\psi\right)=\prod_{j=1}^{k}\varphi\left(t\mid Y_{j},\psi\right)$, respectively. Assumptions $(B4)-(B6)$ now read:
@@ 3. $(B6)$
[…]
Some discussion on the interpretation and applicability of assumptions $(B4)$-$(B6)$ can be found in Appendix B of [3].
Appendix D Proofs
D.1 Proof of Theorem 1
First we introduce some notation and a preliminary lemma. For every $A\subseteq\mathcal{X}$, $\textbf{x}\in\mathcal{X}$ and $i=1,\dots,d$ we write
$A_{i}\left(\textbf{x}_{-i}\right)=A\cap\{\textbf{y}\in\mathcal{X}\,:\ y_{j}=x_{j},j\neq i\},\quad\text{and}\quad A_{i}^{c}\left(\textbf{x}_{-i}\right)=A^{c}\cap\{\textbf{y}\in\mathcal{X}\,:\ y_{j}=x_{j},j\neq i\}\,.$
@@ D.1 Proof of Theorem 1
where, with a slight abuse of notation, $\pi(\text{d}\textbf{x}_{-i})$ denotes the marginal distribution of $X_{-i}$ under $X\sim\pi$. By (8) we have
$\frac{\int_{A_{i}(\textbf{x}_{i})}P_{i}^{\textbf{x}_{-i}}\left(x_{i},A_{i}^{c}(\textbf{x}_{-i})\right)\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})}{\pi_{i}\left(A_{i}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)\pi_{i}\left(A_{i}^{c}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)}\geq\kappa\left(P_{i}^{\textbf{x}_{-i}}\right)$
@@
as desired. If $P_{i}^{\textbf{x}_{-i}}$ is reversible and positive semi-definite, it is possible to show that
@@
Moreover, since $P_{i}^{\textbf{x}_{-i}}$ is positive semi-definite by e.g. Lemma $1.1$ in [39] we have
@@ D.2 Proof of Corollary 1
$P(\partial A)$ $=\frac{1}{d}\sum_{i=1}^{d}P_{i}(\partial A)\geq\frac{1}{d}\sum_{i=1}^{d}\kappa_{i}(P_{i},\mathcal{X})G_{i}(\partial A)\geq\kappa(P.\mathcal{X})\frac{1}{d}\sum_{i=1}^{d}G_{i}(\partial A).$
The first inequality in (9) is obtained dividing by $\pi(A)$ and taking the infimum over $A\subset\mathcal{X}$ such that $1/2\geq\pi(A)>s$. The inequalities in (49) follow by dividing by, respectively, $\pi(A)-s$ or $\pi(A)$ and taking the infimum over $A\subset\mathcal{X}$ such that $1/2\geq\pi(A)>s$.
@@
Finally, if $P^{\textbf{x}_{-i}}$ is reversible and positive semi-definite, again by Theorem 1 we have

### v2

@@ A.2 Convergence of the stationary distributions does not imply convergence of the conductances
Appendix B Reversible and positive semi-definite coordinate-wise Markov chains
Let $P$ as in (5). The next lemma shows that reversibility and positive semi-definiteness of $P$ follow by the ones of $P_{i}$ and in turn of $P_{i}^{\textbf{x}_{-i}}$.
Lemma 5.
Let $P=\frac{1}{d}\sum_{i=1}^{d}P_{i}$ be a Markov kernel. Then, if $P_{i}$ is $\pi$-reversible and positive semi-definite for every $i$, then the same holds for $P$.
Moreover, let $P_{i}$ be as in (5). Then $P_{i}$ is $\pi$-reversible and positive semi-definite if and only if $P_{i}^{\textbf{x}_{-i}}$ is $\pi(\cdot\mid\textbf{x}_{-i})$-reversible and positive semi-definite for $\pi(\text{d}\textbf{x}_{-i})$-almost every $\textbf{x}_{-i}$.
Proof.
If $A,B\subset\mathcal{X}$, then
$\int_{A}\int_{B}P(\textbf{x},\text{d}\textbf{y})\pi(\text{d}\textbf{x})$ $=\frac{1}{d}\sum_{i=1}^{d}\int_{A}\int_{B}P_{i}(\textbf{x},\text{d}\textbf{y})\pi(\text{d}\textbf{x})$
$=\frac{1}{d}\sum_{i=1}^{d}\int_{B}\int_{A}P_{i}(\textbf{x},\text{d}\textbf{y})\pi(\text{d}\textbf{x})=\int_{B}\int_{A}P(\textbf{x},\text{d}\textbf{y})\pi(\text{d}\textbf{x}),$
where the second equality follows by $\pi$-reversibility of $P_{i}$. Moreover, let $f\,:\,\mathcal{X}\,\to\,\mathcal{X}$ be such that $\int f^{2}(\textbf{x})\pi(\text{d}\textbf{x})<\infty$ and notice that
$\int\left(Pf(\textbf{x})\right)f(\textbf{x})\pi(\text{d}\textbf{x})$ $=\frac{1}{d}\sum_{i=1}^{d}\int\left(P_{i}f(\textbf{x})\right)f(\textbf{x})\pi(\text{d}\textbf{x})\geq 0,$
by positive semi-definiteness of $P_{i}$. Thus, $P$ is $\pi$-reversible and positive semi-definite.
As regards the second point, assume that $P_{i}^{\textbf{x}_{-i}}$ is $\pi(\cdot\mid\textbf{x}_{-i})$-reversible and positive semi-definite for $\pi(\text{d}\textbf{x}_{-i})$-almost every $\textbf{x}_{-i}$. Then, letting $A(\textbf{x}_{-i})=\left\{x_{i}\,;\,(x_{i},\textbf{x}_{-i})\in A\right\}$ and similarly for $B(\textbf{x}_{-i})$, we have that
$\int_{A}\int_{B}P_{i}(\textbf{x},\text{d}\textbf{y})\pi(\text{d}\textbf{x})$ $=\int\left[\int_{A(\textbf{x}_{-i})}\int_{B(\textbf{x}_{-i})}P_{i}^{\textbf{x}_{-i}}(x_{i},\text{d}y_{i})\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})\right]\pi(\text{d}\textbf{x}_{-i})$
$=\int\left[\int_{B(\textbf{x}_{-i})}\int_{A(\textbf{x}_{-i})}P_{i}^{\textbf{x}_{-i}}(x_{i},\text{d}y_{i})\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})\right]\pi(\text{d}\textbf{x}_{-i})$
$=\int_{B}\int_{A}P_{i}(\textbf{x},\text{d}\textbf{y})\pi(\text{d}\textbf{x}).$
Moreover
$\int\left(P_{i}f(\textbf{x})\right)f(\textbf{x})\pi(\text{d}\textbf{x})=\int\left[\int\left(P_{i}^{\textbf{x}_{-i}}f(\textbf{x})\right)f(\textbf{x})\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})\right]\pi(\text{d}\textbf{x}_{-i})\geq 0,$
so that $P_{i}$ is $\pi$-reversible and positive semi-definite.
Conversely, assume that $C=\left\{\textbf{x}_{-i}\,;\,P_{i}^{\textbf{x}_{-i}}\text{ not reversible}\right\}\subset\mathcal{X}_{-i}$ is such that $\pi(C)>0$. Then there exist classes $\{A(\textbf{x}_{-i})\}_{\textbf{x}_{-i}\in C}$ and $\{B(\textbf{x}_{-i})\}_{\textbf{x}_{-i}\in C}$ such that $A(\textbf{x}_{-i})\subset\mathcal{X}_{i}$, $B(\textbf{x}_{-i})\subset\mathcal{X}_{i}$ and
$\int_{A(\textbf{x}_{-i})}\int_{B(\textbf{x}_{-i})}P_{i}^{\textbf{x}_{-i}}(x_{i},\text{d}y_{i})\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})>\int_{B(\textbf{x}_{-i})}\int_{A(\textbf{x}_{-i})}P_{i}^{\textbf{x}_{-i}}(x_{i},\text{d}y_{i})\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})$
for every $\textbf{x}_{-i}\in C$. Then, let
$A=\left\{\textbf{x}\,;\,\textbf{x}_{-i}\in C,x_{i}\in A(\textbf{x}_{-i})\right\},\quad\text{and}\quad B=\left\{\textbf{x}\,;\,\textbf{x}_{-i}\in C,x_{i}\in B(\textbf{x}_{-i})\right\}$
and notice that
$\int_{A}\int_{B}P_{i}(\textbf{x},\text{d}\textbf{y})\pi(\text{d}\textbf{x})$ $=\int_{C}\left[\int_{A(\textbf{x}_{-i})}\int_{B(\textbf{x}_{-i})}P_{i}^{\textbf{x}_{-i}}(x_{i},\text{d}y_{i})\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})\right]\pi(\text{d}\textbf{x}_{-i})$
$>\int_{C}\left[\int_{B(\textbf{x}_{-i})}\int_{A(\textbf{x}_{-i})}P_{i}^{\textbf{x}_{-i}}(x_{i},\text{d}y_{i})\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})\right]\pi(\text{d}\textbf{x}_{-i})$
$=\int_{A}\int_{B}P_{i}(\textbf{x},\text{d}\textbf{y})\pi(\text{d}\textbf{x}),$
so that $P_{i}$ cannot be $\pi$-reversible. Similarly, if $C=\left\{\textbf{x}_{-i}\,;\,P_{i}^{\textbf{x}_{-i}}\text{ not positive semi-definite}\right\}\subset\mathcal{X}_{-i}$ is such that $\pi(C)>0$, then there exists a class $\{g^{\textbf{x}_{-i}}\}_{\textbf{x}_{-i}\in C}$ such that $g^{\textbf{x}_{-i}}\,:\,\mathcal{X}_{i}\,\to\,\mathcal{X}_{i}$ and
$\int\left(P_{i}^{\textbf{x}_{-i}}g^{\textbf{x}_{-i}}(x_{i})\right)g^{\textbf{x}_{-i}}(x_{i})\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})<0,$
for every $\textbf{x}_{-i}\in C$. Then, letting $f(\textbf{x})=\mathbbm{1}_{C}(\textbf{x}_{-i})g^{\textbf{x}_{-i}}(x_{i})$ we have that
$\int\left(P_{i}f(\textbf{x})\right)f(\textbf{x})\pi(\text{d}\textbf{x})=\int_{C}\left[\int\left(P_{i}^{\textbf{x}_{-i}}g^{\textbf{x}_{-i}}(x_{i})\right)g^{\textbf{x}_{-i}}(x_{i})\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})\right]\pi(\text{d}\textbf{x}_{-i})<0,$
so that $P_{i}$ cannot be positive semi-definite. ∎
Appendix C Alternative definition of $s$-conductance
@@
Note that $\Phi_{s}(P)\leq\tilde{\Phi}_{s}(P)$ for every $s\geq 0$ and $\tilde{\Phi}_{0}(P)=\Phi_{0}(P)=\Phi(P)$.
@@
Let $P$ be a coordinate-wise operator as in (5), with target distribution $\pi\in\mathcal{P}(\mathcal{X})$, and $\kappa(P,K)$ as in (7). Then for every $0\leq s\leq 1/2$ and $K\subseteq\mathcal{X}$ we have
@@
$P(\partial A)$ $=\frac{1}{d}\sum_{i=1}^{d}P_{i}(\partial A)\geq\frac{1}{d}\sum_{i=1}^{d}\kappa_{i}(P_{i},\mathcal{X})G_{i}(\partial A)\geq\kappa(P,\mathcal{X})\frac{1}{d}\sum_{i=1}^{d}G_{i}(\partial A).$
@@
Appendix D Extension of Corollary 1 to non-uniform updating probabilities
Given probability weights $\textbf{w}=(w_{1},\dots,w_{d})$ such that $w_{i}>0$ and $\sum_{i=1}^{d}w_{i}=1$, we define
$P^{\textbf{w}}=\sum_{i=1}^{d}w_{i}P_{i},\quad P_{i}(\textbf{x},\text{d}\textbf{y})=P_{i}^{\textbf{x}_{-i}}(x_{i},\text{d}y_{i})\delta_{\textbf{x}_{-i}}\left(\text{d}\textbf{y}_{-i}\right).$ (51)
The choice $w_{i}=1/d$ leads to the uniform case in (5). Similarly we define $G^{\textbf{w}}=\sum_{i=1}^{d}w_{i}G_{i}$. The next corollary shows that the same results of Corollary 1 hold for arbitrary weights.
Corollary 3.
Let $P^{\textbf{w}}$ be as in (51). Then for every $s\in[0,1/2)$ and $K\subseteq\mathcal{X}$ we have
$\Phi_{s}(P^{\textbf{w}})$ $\geq\kappa(P^{\textbf{w}},\mathcal{X})\Phi_{s}(G^{\textbf{w}})\quad\text{and}\quad\Phi_{s}(P^{\textbf{w}})\geq\kappa(P^{\textbf{w}},K)\Phi_{s}(G^{\textbf{w}})-\frac{\pi(K^{c})}{s}\left(\sum_{i=1}^{d}w_{i}\kappa_{i}(P_{i},K)\right)\,$, (52)
with $\kappa(P^{\textbf{w}},K)$ as in (7). Moreover, $\Phi_{s}(G^{\textbf{w}})\geq\Phi_{s}(P^{\textbf{w}})$.
Proof.
Let $A\subset\mathcal{X}$. By Theorem 1 we have
$P^{\textbf{w}}(\partial A)$ $=\sum_{i=1}^{d}w_{i}P_{i}(\partial A)\geq\sum_{i=1}^{d}w_{i}\kappa_{i}(P_{i},\mathcal{X})G_{i}(\partial A)\geq\kappa(P^{\textbf{w}},\mathcal{X})\sum_{i=1}^{d}w_{i}G_{i}(\partial A).$
The first inequality in (52) is obtained dividing by $\pi(A)$ and taking the infimum over $A\subset\mathcal{X}$ such that $1/2\geq\pi(A)>s$.
As regards the other inequality, again by Theorem 1 we have
$P^{\textbf{w}}(\partial A)$ $\geq\sum_{i=1}^{d}w_{i}\kappa_{i}(P_{i},K)G_{i}(\partial A)-\left(\sum_{i=1}^{d}w_{i}\kappa_{i}(P_{i},K)\right)\pi(A\cap K^{c})$
$\geq\kappa(P^{\textbf{w}},K)G^{\textbf{w}}(\partial A)-\left(\sum_{i=1}^{d}w_{i}\kappa_{i}(P_{i},K)\right)\pi(K^{c}).$
$\frac{P^{\textbf{w}}(\partial A)}{\pi(A)}\geq\kappa(P^{\textbf{w}},K)\frac{G^{\textbf{w}}(\partial A)}{\pi(A)}-\left(\sum_{i=1}^{d}w_{i}\kappa_{i}(P_{i},K)\right)\frac{\pi(K^{c})}{s},$
from which the right inequality in (52) follows.
Finally, since $P^{\textbf{x}_{-i}}$ is reversible and positive semi-definite, again by Theorem 1 we have
$G^{\textbf{w}}(\partial A)=\sum_{i=1}^{d}w_{i}G_{i}(\partial A)\geq\sum_{i=1}^{d}w_{i}P_{i}(\partial A)=P(\partial A),$
from which we immediately deduce $\Phi_{s}(G^{\textbf{w}})\geq\Phi_{s}(P^{\textbf{w}})$. ∎
Appendix E Regularity assumptions (B4)-(B6) for model (20)
@@
with $s,s^{\prime}=1,\dots S$ and $d=1,\dots,D$. Moreover we write $B_{\delta}$ for the ball of center $\psi^{*}$ and radius $\delta$, and denote expectations with respect to the law of $Y_{j}$ as defined in $(B1)$ by $E_{Y_{j}}[\cdot]$. Finally, we define the posterior characteristic function of $T(\theta_{j})=\left(T_{1}(\theta_{j}),\dots,T_{S}(\theta_{j})\right)$ and $\sum_{j=1}^{k}T(\theta_{j})$, given $\psi$, as $\varphi\left(t\mid Y_{j},\psi\right)=E\left[e^{it^{\top}T(\theta_{j})}\mid Y_{j},\psi\right]$ for $t\in\mathbb{R}^{S}$. and $\varphi^{(k)}\left(t\mid Y_{1:k},\psi\right)=\prod_{j=1}^{k}\varphi\left(t\mid Y_{j},\psi\right)$, respectively. Assumptions (B4)(B6)(B4)-(B6) now read:
@@ 3. $(B6)$
    for almost every $Y_{1},\dots,Y_{k}\overset{\text{iid}}{\sim}Q_{\psi^{*}}$, with $\phi(\epsilon)<1$ for every $\epsilon>0$.
Some discussion on the interpretation and applicability of assumptions (B4)(B4)-$(B6)$ can be found in Appendix B of [4].
Appendix F Proofs
F.1 Proof of Lemma 1
Proof.
Using the notation $\tilde{\Phi}_{s}(P)$ in (48), by Corollary $1.5$ (point (b)) in [45] we have that
$\left\|\mu P^{t}-\pi\right\|_{TV}\leq Ms+M\left(1-\frac{\tilde{\Phi}^{2}_{s}(P)}{2}\right)^{t}\,.$
The result now follows by noticing that $\Phi_{s}(P)\leq\tilde{\Phi}_{s}(P)$ for every $s\geq 0$. ∎
F.2 Proof of Theorem 1
@@
where, with a slight abuse of notation, $\pi(\text{d}\textbf{x}_{-i})$ denotes the marginal distribution of $X_{-i}$ under $X\sim\pi$.
If $\pi_{i}\left(A_{i}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)\leq 1/2$, by (8) we have
$\frac{\int_{A_{i}(\textbf{x}_{i})}P_{i}^{\textbf{x}_{-i}}\left(x_{i},A_{i}^{c}(\textbf{x}_{-i})\right)\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})}{\pi_{i}\left(A_{i}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)\pi_{i}\left(A_{i}^{c}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)}\geq\frac{\int_{A_{i}(\textbf{x}_{i})}P_{i}^{\textbf{x}_{-i}}\left(x_{i},A_{i}^{c}(\textbf{x}_{-i})\right)\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})}{\pi_{i}\left(A_{i}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)}\geq\kappa\left(P_{i}^{\textbf{x}_{-i}}\right).$
If instead $\pi_{i}\left(A_{i}^{c}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)\leq 1/2$, since $P^{\textbf{x}_{-i}}$ is reversible we have
$\frac{\int_{A_{i}(\textbf{x}_{i})}P_{i}^{\textbf{x}_{-i}}\left(x_{i},A_{i}^{c}(\textbf{x}_{-i})\right)\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})}{\pi_{i}\left(A_{i}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)\pi_{i}\left(A_{i}^{c}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)}$ $=\frac{\int_{A_{i}^{c}(\textbf{x}_{i})}P_{i}^{\textbf{x}_{-i}}\left(x_{i},A_{i}(\textbf{x}_{-i})\right)\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})}{\pi_{i}\left(A_{i}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)\pi_{i}\left(A_{i}^{c}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)}$
$\geq\frac{\int_{A_{i}^{c}(\textbf{x}_{i})}P_{i}^{\textbf{x}_{-i}}\left(x_{i},A_{i}(\textbf{x}_{-i})\right)\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})}{\pi_{i}\left(A_{i}^{c}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)}\geq\kappa\left(P_{i}^{\textbf{x}_{-i}}\right).$
In conclusion
$\frac{\int_{A_{i}(\textbf{x}_{i})}P_{i}^{\textbf{x}_{-i}}\left(x_{i},A_{i}^{c}(\textbf{x}_{-i})\right)\pi_{i}(\text{d}x_{i}\mid\textbf{x}_{-i})}{\pi_{i}\left(A_{i}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)\pi_{i}\left(A_{i}^{c}\left(\textbf{x}_{-i}\right)\mid\textbf{x}_{-i}\right)}\geq\kappa\left(P_{i}^{\textbf{x}_{-i}}\right),$
@@
as desired.
As regards the second point, since $P_{i}^{\textbf{x}_{-i}}$ is reversible and positive semi-definite, it is possible to show that
@@
Moreover, since $P_{i}^{\textbf{x}_{-i}}$ is positive semi-definite, by e.g. Lemma $1.1$ in [45] we have
@@ F.3 Proof of Corollary 1
$P(\partial A)$ $=\frac{1}{d}\sum_{i=1}^{d}P_{i}(\partial A)\geq\frac{1}{d}\sum_{i=1}^{d}\kappa_{i}(P_{i},\mathcal{X})G_{i}(\partial A)\geq\kappa(P,\mathcal{X})\frac{1}{d}\sum_{i=1}^{d}G_{i}(\partial A).$
The first inequality in (52) is obtained dividing by $\pi(A)$ and taking the infimum over $A\subset\mathcal{X}$ such that $1/2\geq\pi(A)>s$.
@@
Finally, since $P^{\textbf{x}_{-i}}$ is reversible and positive semi-definite, again by Theorem 1 we have